# Attention layers provably solve single-location regression

### v1

@@ Appendix B Expression of the risk beyond the invariant manifold
$+{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\gamma^{2}-2\gamma^{2}\nu\mathrm{erf}\bigg{(}\frac{\lambda\sqrt{d/2}\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\bigg{)}-2\lambda\gamma^{2}\sqrt{\frac{d}{2}}\eta\theta\frac{1}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\mathrm{erf}^{\prime}\bigg{(}\frac{\lambda\sqrt{d/2}\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\bigg{)}}$
${\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}-\frac{2\lambda^{2}\gamma^{4}\eta\rho}{1+2\lambda^{2}\gamma^{2}}\mathrm{erf}^{\prime\prime}\bigg{(}\frac{\lambda\sqrt{d/2}\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\bigg{)}+(\frac{d}{2}\theta^{2}+\gamma^{2})\zeta\Big{(}\lambda\sqrt{\frac{d}{2}}\,\kappa,\lambda^{2}\gamma^{2}\Big{)}}$
${\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}+\sqrt{\frac{d}{2}}\left(\theta\rho-\frac{\lambda^{2}\gamma^{2}\rho^{2}\kappa}{1+2\lambda^{2}\gamma^{2}}\right)\frac{4\lambda\gamma^{2}}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\mathrm{erf}\bigg{(}\frac{\lambda\sqrt{d/2}\,\kappa}{\sqrt{(1+4\lambda^{2}\gamma^{2})(1+2\lambda^{2}\gamma^{2})}}\bigg{)}\mathrm{erf}^{\prime}\bigg{(}\frac{\lambda\sqrt{d/2}\,\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\bigg{)}}$
${\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}+\frac{4\lambda^{2}\gamma^{4}\rho^{2}}{\sqrt{\pi}\sqrt{1+4\lambda^{2}\gamma^{2}}(1+2\lambda^{2}\gamma^{2})}\mathrm{erf}^{\prime}\bigg{(}-\frac{\lambda\sqrt{d}\kappa}{\sqrt{1+4\lambda^{2}\gamma^{2}}}\bigg{)}}$
$+{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}(L-1)\left[\zeta(0,\lambda^{2})+\frac{8\lambda^{2}}{\pi\sqrt{1+4\lambda^{2}}(1+2\lambda^{2})}\rho^{2}\right]}+{\color[rgb]{0.68,0.36,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.68,0.36,1}\pgfsys@color@cmyk@stroke{0.32}{0.64}{0}{0}\pgfsys@color@cmyk@fill{0.32}{0.64}{0}{0}\frac{4\lambda^{2}}{(1+2\lambda^{2})\pi}(L-1)(L-2)\rho^{2}}$
$+{\color[rgb]{1,0.47,0.62}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.47,0.62}\pgfsys@color@cmyk@stroke{0}{0.53}{0.38}{0}\pgfsys@color@cmyk@fill{0}{0.53}{0.38}{0}\frac{4\lambda(L-1)\rho}{\sqrt{(1+2\lambda^{2})\pi}}\left(\sqrt{\frac{d}{2}}\theta\,\mathrm{erf}\bigg{(}\frac{\lambda\sqrt{d/2}\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\bigg{)}+\frac{\lambda\gamma^{2}\rho}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\,\mathrm{erf}^{\prime}\bigg{(}\frac{\lambda\sqrt{d/2}\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\right)}\,.$
@@
Appendix C Technical results
@@
Appendix D Experimental details and additional results
Our code is available at
@@ Appendix D Experimental details and additional results
D.1. Experiment of Section 2 (NLP motivations)
Data generation.  We use synthetically-generated data for this experiment. To create our train set, we generate sentences according to the patterns
@@ D.1. Experiment of Section 2 (NLP motivations)
Finally, the fourth test set combines a different sentence pattern and unseen adjectives. The size of the datasets is given in Table 1. All datasets have the same number of $+1$ and $-1$ labels.
@@
Table 1. Size of the generated datasets.
@@
Model.  We recall that there exists several families of Transformer architectures, which in particular are not all best suited for sequence classification. An appropriate family is called encoder-only Transformer, and a foremost example is BERT (Devlin et al., 2019). We refer to Phuong and Hutter (2022) for an introductory discussion of Transformer architectures and associated algorithms. Here, we use a pretrained BERT model from the Hugging Face Transformers library (Wolf et al., 2020), with the default configuration, namely bert-base-uncased. The model has $110$M parameters, $12$ layers, the tokens have dimension $d=768$, and each attention layer has $12$ heads. It was pretrained by masked language modeling, namely some tokens in the input are hidden, and the model learns to predict the missing tokens. We refer to Devlin et al. (2019) for details on the architecture and pretraining procedure. We do not perform any fine-tuning on the model.
@@
Experiment design.  Our experiment consists in performing logistic regression on embeddings of [CLS] tokens in the hidden layers of the pretrained BERT model, where we recall that the [CLS] token is a special token added to the beginning of each input sequence. This is a particular case of the so-called linear probing, which is a common technique in the field of LLMs interpretability. More precisely, let $\ell$ denote a layer index between 00 and $12$, where the index 00 corresponds to the input to the model (after tokenization and embedding in ${\mathbb{R}}^{d}$). Then, for each value of $\ell\in\{0,\dots,12\}$, we train a logistic regression classifier, where, for each example, the input to the classifier is the embedding of the [CLS] token at layer $\ell$ (that is, a $d$-dimensional vector), and the label is simply the label of the sentence as described above.
@@
Results.  For $\ell=0$ (blue bar in Figure 1(b)), the embedding of [CLS] is a fixed vector that does not depend on the rest of the sequence, so the classifier has a pure-chance accuracy of $50\%$. However, as soon as $\ell>0$, thanks to the attention mechanism, the [CLS] token contains information about the sequence. We report in Figure 1(b) the average accuracy over $\ell\in\{1,\dots,12\}$ for the train set (in orange) and the test sets (in green). We observe that the information contained in the [CLS] token is actually very rich, since logistic regression achieves a perfect accuracy of $100\%$ in the train set. In other words, the data fed to the classifier is linearly separable. We emphasize that the size of the train set is significantly larger than the ambient dimension $d$, so it is far from trivial that this procedure would yield a linearly-separable dataset. Therefore, obtaining linearly-separable data demonstrates that the model constructs a linear representations of the input inside the [CLS] token. Moving on to the test sets, the accuracy on the baseline test set is very good ($95\%$), which suggests some generalization abilities of the model. The accuracy on the out-of-distribution test sets degrades (between $64\%$ and $75\%$), but remains largely superior to pure-chance performance. This suggests that the internal representation built by the Transformer model is to some extent universal, in the sense that it is robust to the specifics of the sentence structure and of the word choice.
D.2. Experiment of Section 5 (Gradient descent recovers the oracle predictor)
@@
PGD with an initialization on the sphere and constant inverse temperature schedule.  As emphasized in Section 5, the dynamics of PGD with a general initialization on $(\mathbb{S}^{d-1})^{2}$ depend on the choice of the inverse temperature schedule $\lambda_{t}$. The experiment presented in the main text in Figure 4(a) is for a decreasing schedule $\lambda_{t}=1/(1+10^{-4}t)$. We report in Figure 7 results when taking a constant inverse temperature. We observe distinct patterns depending on the value of this parameter. With a large inverse temperature (Figure 7(a)), we observe that the dynamics in $(\kappa,\nu)$ always escape the neighborhood of 00. Furthermore, the direction $v^{\star}$ is almost perfectly recovered, i.e., $\nu\approx 1$. However, the value of $k^{\star}$ is only partially recovered: the dynamics stabilize around $\kappa\approx 0.3$. Moreover, the excess risk plateaus at a high value, while the dynamics stay far away from the manifold ${\mathcal{M}}$. In the case of a smaller inverse temperature (Figure 7(b)), the situation is different. We observe that some initializations lead to a convergence to the point $(\kappa,\nu)=(0,0)$, in which case the dynamics stay far from the manifold ${\mathcal{M}}$. In other words, there is no recovery of $k^{\star}$ and $v^{\star}$. Other initializations lead to perfect recovery of $k^{\star}$ and $v^{\star}$. In all cases, the final excess risk is low. Theoretical study of these observations is left for future work.
@@
Figure 7. Dynamics of PGD from a random initialization on $(\mathbb{S}^{d-1})^{2}$, for two iteration-independent values of $\lambda_{t}$. Left: Excess risk as a function of the number of steps. Middle left: Alignment $|\kappa|=|k^{\top}k^{\star}|$ and $|\nu|=|v^{\top}v^{\star}|$ with the oracle parameters. Middle right: Trajectories of $\kappa$ and $\nu$ in a few repetitions of the experiments. Each repetition corresponds to a color, the end point of each trajectory is in blue. Right: Distance to the invariant manifold ${\mathcal{M}}$. In all plots except the middle right ones, the experiment is repeated $30$ times with independent random initializations, and either $95\%$ percentile intervals are plotted or all the curves are plotted. Parameters are $d=400$, $L=10$, and $\gamma=\sqrt{1/2}$.
Implementation details.  The implementation of the PGD algorithm (10) requires to compute the gradient of the risk. To this aim, we use the formula for the risk given by Proposition 15. Note that all quantities appearing in this expression have explicit derivatives. The only quantity for which this is not directly clear is the function $\zeta$, which needs to be differentiated with respect to its first variable to compute the derivative of the risk with respect to $\kappa$. However, recall that $\zeta(t,\gamma^{2}):=\mathbb{E}\left[\mathrm{erf}^{2}(t+G)\right]$. Then, by Lemma (18),
@@
Parameter values.  The following table summarizes the value of the parameters in our experiments.

### v2

@@ Appendix C Expression of the risk beyond the invariant manifold
$+{\color[rgb]{.255,.41,.884}\definecolor[named]{pgfstrokecolor}{rgb}{.255,.41,.884}\gamma^{2}-2\gamma^{2}\nu\mathrm{erf}\bigg{(}\frac{\lambda\sqrt{d/2}\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\bigg{)}-2\lambda\gamma^{2}\sqrt{\frac{d}{2}}\eta\theta\frac{1}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\mathrm{erf}^{\prime}\bigg{(}\frac{\lambda\sqrt{d/2}\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\bigg{)}}$
${\color[rgb]{.255,.41,.884}\definecolor[named]{pgfstrokecolor}{rgb}{.255,.41,.884}-\frac{2\lambda^{2}\gamma^{4}\eta\rho}{1+2\lambda^{2}\gamma^{2}}\mathrm{erf}^{\prime\prime}\bigg{(}\frac{\lambda\sqrt{d/2}\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\bigg{)}+(\frac{d}{2}\theta^{2}+\gamma^{2})\zeta\Big{(}\lambda\sqrt{\frac{d}{2}}\,\kappa,\lambda^{2}\gamma^{2}\Big{)}}$
${\color[rgb]{.255,.41,.884}\definecolor[named]{pgfstrokecolor}{rgb}{.255,.41,.884}+\sqrt{\frac{d}{2}}\left(\theta\rho-\frac{\lambda^{2}\gamma^{2}\rho^{2}\kappa}{1+2\lambda^{2}\gamma^{2}}\right)\frac{4\lambda\gamma^{2}}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\mathrm{erf}\bigg{(}\frac{\lambda\sqrt{d/2}\,\kappa}{\sqrt{(1+4\lambda^{2}\gamma^{2})(1+2\lambda^{2}\gamma^{2})}}\bigg{)}\mathrm{erf}^{\prime}\bigg{(}\frac{\lambda\sqrt{d/2}\,\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\bigg{)}}$
${\color[rgb]{.255,.41,.884}\definecolor[named]{pgfstrokecolor}{rgb}{.255,.41,.884}+\frac{4\lambda^{2}\gamma^{4}\rho^{2}}{\sqrt{\pi}\sqrt{1+4\lambda^{2}\gamma^{2}}(1+2\lambda^{2}\gamma^{2})}\mathrm{erf}^{\prime}\bigg{(}-\frac{\lambda\sqrt{d}\kappa}{\sqrt{1+4\lambda^{2}\gamma^{2}}}\bigg{)}}$
$+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(L-1)\left[\zeta(0,\lambda^{2})+\frac{8\lambda^{2}}{\pi\sqrt{1+4\lambda^{2}}(1+2\lambda^{2})}\rho^{2}\right]}+{\color[rgb]{.73,.332,.828}\definecolor[named]{pgfstrokecolor}{rgb}{.73,.332,.828}\frac{4\lambda^{2}}{(1+2\lambda^{2})\pi}(L-1)(L-2)\rho^{2}}$
$+{\color[rgb]{.98,.5,.448}\definecolor[named]{pgfstrokecolor}{rgb}{.98,.5,.448}\frac{4\lambda(L-1)\rho}{\sqrt{(1+2\lambda^{2})\pi}}\left(\sqrt{\frac{d}{2}}\theta\,\mathrm{erf}\bigg{(}\frac{\lambda\sqrt{d/2}\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\bigg{)}+\frac{\lambda\gamma^{2}\rho}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\,\mathrm{erf}^{\prime}\bigg{(}\frac{\lambda\sqrt{d/2}\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\right)}\,.$
@@
Appendix D Technical results
@@
This lemma reveals the importance of choosing the $\mathrm{erf}$ function as the component-wise nonlinearity: there are closed-form formulas for the expectation of $\mathrm{erf}$ and its derivatives applied to Gaussian random variables. Extending the results to any nonlinear, bounded, increasing, equal to 0 at 0, and differentiable activation function is an interesting next step.
@@
Appendix E Experimental details and additional results
@@
E.1 Experiment of Section 2 (NLP motivations)
Data generation.
We use synthetically-generated data for this experiment. To create our train set, we generate sentences according to the patterns
@@ E.1 Experiment of Section 2 (NLP motivations)
Finally, the fourth test set combines a different sentence pattern and unseen adjectives. The size of the datasets is given in the table below. All datasets have the same number of $+1$ and $-1$ labels.
@@
Table 1: Size of the generated datasets.
Model.
@@
We recall that there exists several families of Transformer architectures, which in particular are not all best suited for sequence classification. An appropriate family is called encoder-only Transformer, and a foremost example is BERT (Devlin et al., 2019). We refer to Phuong & Hutter (2022) for an introductory discussion of Transformer architectures and associated algorithms. Here, we use a pretrained BERT model from the Hugging Face Transformers library (Wolf et al., 2020), with the default configuration, namely bert-base-uncased. The model has $110$M parameters, $12$ layers, the tokens have dimension $d=768$, and each attention layer has $12$ heads. It was pretrained by masked language modeling, namely some tokens in the input are hidden, and the model learns to predict the missing tokens. We refer to Devlin et al. (2019) for details on the architecture and pretraining procedure. We do not perform any fine-tuning on the model.
Experiment design.
@@
Our experiment consists in performing logistic regression on embeddings of [CLS] tokens in the hidden layers of the pretrained BERT model, where we recall that the [CLS] token is a special token added to the beginning of each input sequence. This is a particular case of the so-called linear probing, which is a common technique in the field of LLMs interpretability. More precisely, let $\ell$ denote a layer index between 00 and $12$, where the index 00 corresponds to the input to the model (after tokenization and embedding in ${\mathbb{R}}^{d}$). Then, for each value of $\ell\in\{0,\dots,12\}$, we train a logistic regression classifier, where, for each example, the input to the classifier is the embedding of the [CLS] token at layer $\ell$ (that is, a $d$-dimensional vector), and the label is simply the label of the sentence as described above.
Results.
@@
For $\ell=0$ (blue bar in Figure 1(b)), the embedding of [CLS] is a fixed vector that does not depend on the rest of the sequence, so the classifier has a pure-chance accuracy of $50\%$. However, as soon as $\ell>0$, thanks to the attention mechanism, the [CLS] token contains information about the sequence. We report in Figure 1(b) the average accuracy over $\ell\in\{1,\dots,12\}$ for the train set (in orange) and the test sets (in green). We observe that the information contained in the [CLS] token is actually very rich, since logistic regression achieves a perfect accuracy of $100\%$ in the train set. In other words, the data fed to the classifier is linearly separable. We emphasize that the size of the train set is significantly larger than the ambient dimension $d$, so it is far from trivial that this procedure would yield a linearly-separable dataset. Therefore, obtaining linearly-separable data demonstrates that the model constructs a linear representations of the input inside the [CLS] token. Moving on to the test sets, the accuracy on the baseline test set is very good ($95\%$), which suggests some generalization abilities of the model. The accuracy on the out-of-distribution test sets degrades (between $64\%$ and $75\%$), but remains largely superior to pure-chance performance. This suggests that the internal representation built by the Transformer model is to some extent universal, in the sense that it is robust to the specifics of the sentence structure and of the word choice.
E.2 Experiment of Section 5 (Gradient descent recovers the oracle predictor)
@@
PGD with an initialization on the sphere and constant inverse temperature schedule.
As emphasized in Section 5, the dynamics of PGD with a general initialization on $(\mathbb{S}^{d-1})^{2}$ depend on the choice of the inverse temperature schedule $\lambda_{t}$. The experiment presented in the main text in Figure 4(a) is for a decreasing schedule $\lambda_{t}=1/(1+10^{-4}t)$. We report in Figure 7 results when taking a constant inverse temperature. We observe distinct patterns depending on the value of this parameter. With a large inverse temperature (Figure 7(a)), we observe that the dynamics in $(\kappa,\nu)$ always escape the neighborhood of 00. Furthermore, the direction $v^{\star}$ is almost perfectly recovered, i.e., $\nu\approx 1$. However, the value of $k^{\star}$ is only partially recovered: the dynamics stabilize around $\kappa\approx 0.3$. Moreover, the excess risk plateaus at a high value, while the dynamics stay far away from the manifold ${\mathcal{M}}$. In the case of a smaller inverse temperature (Figure 7(b)), the situation is different. We observe that some initializations lead to a convergence to the point $(\kappa,\nu)=(0,0)$, in which case the dynamics stay far from the manifold ${\mathcal{M}}$. In other words, there is no recovery of $k^{\star}$ and $v^{\star}$. Other initializations lead to perfect recovery of $k^{\star}$ and $v^{\star}$. In all cases, the final excess risk is low. Theoretical study of these observations is left for future work.
@@
Figure 7: Dynamics of PGD from a random initialization on $(\mathbb{S}^{d-1})^{2}$, for two iteration-independent values of $\lambda_{t}$. Left: Excess risk as a function of the number of steps. Middle left: Alignment $|\kappa|=|k^{\top}k^{\star}|$ and $|\nu|=|v^{\top}v^{\star}|$ with the oracle parameters. Middle right: Trajectories of $\kappa$ and $\nu$ in a few repetitions of the experiments. Each repetition corresponds to a color, the end point of each trajectory is in blue. Right: Distance to the invariant manifold ${\mathcal{M}}$. In all plots except the middle right ones, the experiment is repeated $30$ times with independent random initializations, and either $95\%$ percentile intervals are plotted or all the curves are plotted. Parameters are $d=400$, $L=10$, and $\gamma=\sqrt{1/2}$.
Implementation details.
The implementation of the PGD algorithm (10) requires to compute the gradient of the risk. To this aim, we use the formula for the risk given by Proposition 15. Note that all quantities appearing in this expression have explicit derivatives. The only quantity for which this is not directly clear is the function $\zeta$, which needs to be differentiated with respect to its first variable to compute the derivative of the risk with respect to $\kappa$. However, recall that $\zeta(t,\gamma^{2}):=\mathbb{E}\left[\mathrm{erf}^{2}(t+G)\right]$. Then, by Lemma (18),
@@
Parameter values.
The following table summarizes the value of the parameters in our experiments.